OpenStudy (anonymous):

The graph of f(x)=1/x^2-c has a vertical asymptote at x=3. Find c. PreCalculus help please

OpenStudy (anonymous):

find c?

OpenStudy (anonymous):

tell me if 1/x^2-c is a fraction form?

OpenStudy (anonymous):

Yes it is in fraction form, 1 on top and x^2-c on the bottom

OpenStudy (anonymous):

this is calculus y do u want to find c

OpenStudy (anonymous):

I don't know... It confuses me... This is more of an algebra question than calculus.. and I sucked at algebra....

OpenStudy (anonymous):

i cant find c with f(x) thingie

OpenStudy (anonymous):

Okay.. Thanks anyway :)

OpenStudy (anonymous):

c=x^2-1/y there u go

myininaya (myininaya):

If f has a vertical asymptote at x=3, then that means the function shouldn't exist at x=3. The function doesn't exist when the fraction doesn't exist. The fraction doesn't exist when the denominator is 0 (when the bottom is 0). When is x^2-c equal to 0? And since the vertical asymptote is at x=3 then we no 3^2-c equal to 0 will give us a value of c that makes the function undefined at x=3.

myininaya (myininaya):

know not no (darn it)

OpenStudy (anonymous):

so c=9?

myininaya (myininaya):

Yep! That would give us f(x)=1/(x^2-9) As you can see this function does not exist at x=3 (or x=-3 <---but whatever :p)

OpenStudy (anonymous):

Awesome!! Thanks a million!!! Do you think you could help me with this one too? Find the equation of the slant asymptote for the graph of y=\[y=\frac{ 3x ^{2}+2x-3 }{ x-1 }\] @myininaya

myininaya (myininaya):

All you really need to do is divide.

OpenStudy (anonymous):

I know I have to turn it into a division problem.. But after that I'm not sure... Because the long division with Xs confuses me..

myininaya (myininaya):

If you can do the division correctly and if the numerator is 1 degree higher than the numerator, then the slant asymptote will be the quotient.

myininaya (myininaya):